a 49.8 kg crate is at rest on a level floor, and the coefficient of kinetic friction is 0.327. The acceleration of gravity is 9.8 m/s^2. If the crate is pushed horizontally with a force of 231.404 N, how far does it move in 7.33 s?
1
Expert's answer
2012-10-18T11:31:12-0400
Fgr=mg
Ffr=kFgr=159.589 N

F-Fr=71.815 N

F=ma

a=F/m=1.44 m/s^2

at^2/2=S

S=38.74 m